An electric car is being driven through a town at 22 m/s. The kinetic energy of the car is 222640J. Calculate the mass of the car


1
Expert's answer
2021-01-19T07:07:49-0500

By the definition of the kinetic energy, we have:


"KE=\dfrac{1}{2}mv^2,""m=\dfrac{2KE}{v^2},""m=\dfrac{2\cdot222640\ J}{(22\ \dfrac{m}{s})^2}=920\ kg."

Answer:

"m=920\ kg."